OpenStudy (anonymous):

Points Y and Z are the same distance from point X on a number line. The coordinate of point X is 8 and the coordinate of point Y is 2. What is the coordinate of point Z?

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

Hello

OpenStudy (p0sitr0n):

z=14 since x-y=6, z=x+6=14

OpenStudy (princeharryyy):



OpenStudy (princeharryyy):

@XxAminaxX

OpenStudy (princeharryyy):

ans(16,4)

OpenStudy (princeharryyy):

did u get it @XxAminaxX

OpenStudy (anonymous):

@princeharryyy @POstir0n Yeah the correct answer was 14

OpenStudy (princeharryyy):

ohk may b I got it wrong somewhere @XxAminaxX

OpenStudy (princeharryyy):

actually I thought that the centre point was given as(8,2) what i wrote in my figure but actually they have given something else... Sorry, I read it wrong @XxAminaxX

OpenStudy (princeharryyy):

So, yeah the correct answer would be 14 not 16. Sorry for giving the wrong answer... I interpreted it wrong...

OpenStudy (anonymous):

It's okay, doesn't really matter anymore :P @princeharryyy

OpenStudy (princeharryyy):

whenever u tell someone the wrong may be the reason could be any.... u should apologise ... if u r wrong... no matter what the consequences would b...) @XxAminaxX I kno mistakes happen but u should apologise if u do something wrong...

OpenStudy (anonymous):

Not if it didn't effect me, I kinda already knew the answer but I needed to check if I was right @princeharryyy

OpenStudy (princeharryyy):

it does not matter .... here but there may b circumstances where it matters alot... @XxAminaxX ... well, it was nice if it didn't effect u.... :) have a nice day!

OpenStudy (anonymous):

You too, bye! @princeharryyy

OpenStudy (princeharryyy):

bieee :p @XxAminaxX